OpenStudy (thatonekidkobe):

Which of the expressions below can be factored using the difference of squares method? A. 17x^2 + 23y^2 B. 25x^2 - 64y^2 C. 17x^2 - 23y^2 D. 25x^2 + 64y^2

rishavraj (rishavraj):

\[a^2 - b^2 = (a + b)(a - b)\]

OpenStudy (thatonekidkobe):

So B or C because of the subtraction

rishavraj (rishavraj):

@thatonekidkobe wht do u think ??? u see 25 = 5^2 64 = 8^2

OpenStudy (thatonekidkobe):

Oh okay, B it is thanks :)